# Equilibrium Investment with Random Risk Aversion: (Non-)uniqueness, Optimality, and Comparative Statics

### v2

@@ Appendix A Properties of the function $h$
By Hölder’s inequality,
$\left(\tilde{\mathbb{E}}\left[\bm{R}e^{-\frac{1}{2}\bm{R}x}\right]\right)^{2}=\left(\tilde{\mathbb{E}}\left[\bm{R}e^{-\frac{1}{2}\bm{R}x}\right]\cdot\tilde{\mathbb{E}}\left[\bm{R}e^{-\frac{1}{2}\bm{R}x}\right]\right)\leq\tilde{\mathbb{E}}\left[e^{-\frac{1}{2}\bm{R}x}\right]\tilde{\mathbb{E}}\left[\bm{R}^{2}e^{-\frac{1}{2}\bm{R}x}\right].$ (A.1)
Hence, $h^{\prime}\geq 0$ and $h$ is increasing. The inequality in (A.1) becomes an equality if and only if $\tilde{\mathbb{E}}\left[e^{-\frac{1}{2}\bm{R}x}\right]\bm{R}^{2}e^{-\frac{1}{2}\bm{R}x}=e^{-\frac{1}{2}\bm{R}x}\tilde{\mathbb{E}}\left[\bm{R}^{2}e^{-\frac{1}{2}\bm{R}x}\right]$, which is equivalent to $\bm{R}$ following a single-point distribution. Therefore, $h$ is strictly increasing unless $\bm{R}$ is a constant. ∎
@@
Let $\bm{R}$ be a non-negative random variable, and let $r_{0}\triangleq\operatorname{essinf}\bm{R}$. Then,
$\lim_{x\to\infty}h(x)=\frac{1}{r_{0}},$
where we adopt the convention that $\frac{1}{0}=\infty$.
@@
where $Q(x)\triangleq\frac{\tilde{\mathbb{E}}[\bm{X}e^{-\frac{1}{2}\bm{X}x}]}{\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{X}x}]}$. For any arbitrary $\varepsilon>0$, because $\operatorname{essinf}\bm{X}=0$, there exists a probability mass $p_{\varepsilon}\triangleq\tilde{\mathbb{P}}(0\leq\bm{X}\leq\varepsilon)>0$. Therefore,
$Q(x)\leq\frac{\tilde{\mathbb{E}}[\bm{X}e^{-\frac{1}{2}\bm{X}x}\mathbb{I}_{\{\bm{X}\leq\varepsilon\}}+\tilde{\mathbb{E}}[\bm{X}e^{-\frac{1}{2}\bm{X}x}\mathbb{I}_{\{\bm{X}>\varepsilon\}}]}{\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{X}x}\mathbb{I}_{\{\bm{X}\leq\varepsilon\}}]}\leq\varepsilon+\frac{\tilde{\mathbb{E}}[\bm{X}e^{-\frac{1}{2}\bm{X}x}\mathbb{I}_{\{\bm{X}>\varepsilon\}}]}{p_{\varepsilon}e^{-\frac{1}{2}\varepsilon x}}=\varepsilon+\frac{\tilde{\mathbb{E}}[\bm{X}e^{-\frac{1}{2}(\bm{X}-\varepsilon)x}\mathbb{I}_{\{\bm{X}>\varepsilon\}}]}{p_{\varepsilon}}.$
Note that for any $x\geq 1$, $\bm{X}e^{-\frac{1}{2}(\bm{X}-\varepsilon)x}\mathbb{I}_{\{\bm{X}>\varepsilon\}}$ is dominated by $\bm{X}e^{-\frac{1}{2}(\bm{X}-\varepsilon)}\mathbb{I}_{\{\bm{X}>\varepsilon\}}$, which is integrable regardless of whether $\tilde{\mathbb{E}}[\bm{X}]$ is finite. Hence, by the dominated convergence theorem, $\limsup\limits_{x\to\infty}Q(x)\leq\varepsilon$. Because $\varepsilon$ is arbitrary, we have $\lim\limits_{x\to\infty}Q(x)=0$. By (A.2), the convergence result for $h$ follows immediately. ∎
@@
Because $\tilde{\mathbb{E}}[\bm{R}_{n}]\to\tilde{\mathbb{E}}[\bm{R}]$, the sequence $\{\tilde{\mathbb{E}}[\bm{R}_{n}]\}_{n\geq 1}$ is bounded. Hence, the family of functions $\{\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{R}_{n}x}]\}_{n\geq 1}$ is equicontinuous on $\mathcal{K}$. Moreover, because $\bm{R}_{n}\xrightarrow{d}\bm{R}$ and the function $r\mapsto e^{-\frac{1}{2}rx}$ is bounded and continuous, we have $\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{R}_{n}x}]\to\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{R}x}]$ for any $x\geq 0$. It follows that $\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{R}_{n}x}]\to\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{R}x}]$ uniformly on $\mathcal{K}$; see, e.g., rudin1976principles.
Now we show the uniform convergence of $\{\tilde{\mathbb{E}}[\bm{R}_{n}e^{-\frac{1}{2}\bm{R}_{n}x}]\}_{n\geq 1}$ on $\mathcal{K}$. By $\bm{R}_{n}\xrightarrow{d}\bm{R}$ and $\tilde{\mathbb{E}}[\bm{R}_{n}]\to\tilde{\mathbb{E}}[\bm{R}]<\infty$, we know that the family $\{\bm{R}_{n}\}_{n\geq 1}$ is uniformly integrable. Consequently, for any $\varepsilon>0$, there exists $M>0$ such that $F_{\bm{R}}$ is continuous at $M$, $\sup_{n}\tilde{\mathbb{E}}[\bm{R}_{n}\mathbb{I}_{\{\bm{R}_{n}>M\}}]<\frac{\varepsilon}{3}$, and $\tilde{\mathbb{E}}[\bm{R}\mathbb{I}_{\{\bm{R}>M\}}]<\frac{\varepsilon}{3}$. Let $f_{n}^{M}(x)\triangleq\tilde{\mathbb{E}}[\bm{R}_{n}e^{-\frac{1}{2}\bm{R}_{n}x}\mathbb{I}_{\{\bm{R}_{n}\leq M\}}]$. Then $f_{n}^{M}$ is continuously differentiable and its derivative is uniformly bounded by $\frac{1}{2}M^{2}$, which implies that the family $\{f_{n}^{M}\}_{n\geq 1}$ is equicontinuous on $\mathcal{K}$. Furthermore, because $\bm{R}_{n}\xrightarrow{d}\bm{R}$, we have the pointwise convergence $f_{n}^{M}(x)\to f^{M}(x)\triangleq\tilde{\mathbb{E}}[\bm{R}e^{-\frac{1}{2}\bm{R}x}\mathbb{I}_{\{\bm{R}\leq M\}}]$; see, e.g., follmer2016stochastic. As a result, $f_{n}^{M}$ converges uniformly to $f^{M}$ on $\mathcal{K}$. Hence, for all sufficiently large $n$,
@@
$|\tilde{\mathbb{E}}[\bm{R}_{n}e^{-\frac{1}{2}\bm{R}_{n}x}]-\tilde{\mathbb{E}}[\bm{R}e^{-\frac{1}{2}\bm{R}x}]|\leq|f_{n}^{M}(x)-f^{M}(x)|+\tilde{\mathbb{E}}[\bm{R}_{n}\mathbb{I}_{\{\bm{R}_{n}>M\}}]+\tilde{\mathbb{E}}[\bm{R}\mathbb{I}_{\{\bm{R}>M\}}].$
@@
Finally, because $\tilde{\mathbb{E}}[\bm{R}]>0$, the function $\tilde{\mathbb{E}}[\bm{R}e^{-\frac{1}{2}\bm{R}x}]$ is bounded away from zero on $\mathcal{K}$. Therefore, $\{h_{\bm{R}_{n}}\}_{n\geq 1}$ converges uniformly to $h_{\bm{R}}$ on $\mathcal{K}$. ∎
The reverse hazard rate order can be characterized via expectation ratios and connected to first-order stochastic dominance under weighted measures, as established by caperaa1988tail (see also SS2007) and wang2024weighted.
Lemma A.4 (caperaa1988tail).
Let $\bm{R_{1}}$ and $\bm{R_{2}}$ be two random variables. Then $R_{1}\succeq_{rh}R_{2}$ if and only if
$\frac{\tilde{\mathbb{E}}\left[u(\bm{R_{1}})w(\bm{R_{1}})\right]}{\tilde{\mathbb{E}}[w(\bm{R_{1}})]}\geq\frac{\tilde{\mathbb{E}}[u(\bm{R_{2}})w(\bm{R_{2}})]}{\tilde{\mathbb{E}}[w(\bm{R_{2}})]},$
for all functions $u$ and $w$ for which the expectations exist, such that $w$ is positive, continuous, and decreasing, and $u$ is continuous and strictly increasing.
Lemma A.5 (wang2024weighted).
Let $\bm{R_{1}}$ and $\bm{R_{2}}$ be two random variables with cumulative distribution functions $F_{\bm{R_{1}}}$ and $F_{\bm{R_{2}}}$, respectively. Assume that $\bm{R_{1}}\succeq_{rh}\bm{R_{2}}$. Let $w(\cdot)$ be a positive, continuous, and strictly decreasing function. Define the weighted probability measures $\tilde{F}_{i}$ associated with $\bm{R}_{i}$ for $i=1,2$ by
$\mathrm{d}\tilde{F}_{i}(x)=\frac{w(x)\mathrm{d}F_{\bm{R_{i}}}}{\tilde{\mathbb{E}}\left[w(\bm{R_{i}})\right]}.$ (A.3)
Then, $\bm{R}_{1}$ dominates $\bm{R}_{2}$ under the weighted measures in the sense of first-order stochastic dominance, i.e.,
$\tilde{F}_{1}\succeq_{1}\tilde{F}_{2}.$
[…]
@@
For fixed $x>0$, $h(x)$ is strictly decreasing with respect to $\bm{R}$ in the sense of the reverse-hazard-rate order among distinct distributions.
@@
Fix $x>0$. Let $\bm{R_{1}}$ and $\bm{R_{2}}$} be two non-negative random variables such that $\bm{R_{1}}\succeq_{rh}\bm{R_{2}}$ and their distributions are distinct. For notational simplicity, in the sequel, let $h_{i}(\cdot)$ and $H_{i}(\cdot)$ denote the function $h(\cdot)$ and the function defined in Theorem 3.2 corresponding to the random variable $\bm{R}_{i}$ for $i=1,2$, respectively. According to Lemma A.4, let $u(\gamma)=\gamma$ and $w(\gamma)=\exp(-\frac{1}{2}\gamma x)$. Note that $u(\gamma)$ is a strictly increasing function, and for $x>0$, $w(\gamma)$ is a positive, continuous, and strictly decreasing function of $\gamma$. Substituting them into this inequality, we have
$\frac{\tilde{\mathbb{E}}[\bm{R_{1}}\exp(-\frac{1}{2}\bm{R_{1}}x)]}{\tilde{\mathbb{E}}[\exp(-\frac{1}{2}\bm{R_{1}}x)]}\geq\frac{\tilde{\mathbb{E}}[\bm{R_{2}}\exp(-\frac{1}{2}\bm{R_{2}}x)]}{\tilde{\mathbb{E}}[\exp(-\frac{1}{2}\bm{R_{2}}x)]}.$
Moreover, assume that the equality holds. Define the weighted probability measures $\tilde{F}_{i}$ for $i=1,2$ by (A.3). The given condition is equivalent to $\tilde{\mathbb{E}}_{\tilde{F}_{1}}[\bm{R_{1}}]=\tilde{\mathbb{E}}_{\tilde{F}_{2}}[\bm{R_{1}}]$. From Lemma A.5, we have $\tilde{F}_{1}\succeq_{1}\tilde{F}_{2}$ and consequently, $\mathrm{d}\tilde{F}_{1}(\gamma)=\mathrm{d}\tilde{F}_{2}(\gamma)$. Given that $w(\gamma)$ is positive, this implies $\frac{\mathrm{d}\tilde{F}_{1}}{\tilde{\mathbb{E}}\left[w(\bm{R_{1}})\right]}=\frac{\mathrm{d}\tilde{F}_{2}}{\tilde{\mathbb{E}}\left[w(\bm{R_{2}})\right]}$. Integrating both sides over the entire support and using the property $\int\mathrm{d}\tilde{F}_{i}(\gamma)=1$, we obtain $1=\frac{\tilde{\mathbb{E}}\left[w(\bm{R_{1}})\right]}{\tilde{\mathbb{E}}\left[w(\bm{R_{2}})\right]}$, which implies $\tilde{\mathbb{E}}\left[w(\bm{R_{1}})\right]=\tilde{\mathbb{E}}\left[w(\bm{R_{2}})\right]$. Substituting this back yields $\mathrm{d}F_{\bm{R_{1}}}(\gamma)=\mathrm{d}F_{\bm{R_{2}}}(\gamma)$, confirming that $F_{\bm{R_{1}}}(\gamma)=F_{\bm{R_{2}}}(\gamma)$. This contradicts the different distributions of $\bm{R_{1}}$ and $\bm{R_{2}}$. Thus, we have
$\frac{\tilde{\mathbb{E}}[\bm{R_{1}}\exp(-\frac{1}{2}\bm{R_{1}}x)]}{\tilde{\mathbb{E}}[\exp(-\frac{1}{2}\bm{R_{1}}x)]}>\frac{\tilde{\mathbb{E}}[\bm{R_{2}}\exp(-\frac{1}{2}\bm{R_{2}}x)]}{\tilde{\mathbb{E}}[\exp(-\frac{1}{2}\bm{R_{2}}x)]},$
and equivalently, $h_{1}(x)<h_{2}(x)$. ∎

### v3

@@ Appendix A Properties of the function $h$
By the Cauchy-Schwartz inequality,
$\left(\tilde{\mathbb{E}}\left[\bm{R}e^{-\frac{1}{2}\bm{R}x}\right]\right)^{2}\leq\tilde{\mathbb{E}}\left[e^{-\frac{1}{2}\bm{R}x}\right]\tilde{\mathbb{E}}\left[\bm{R}^{2}e^{-\frac{1}{2}\bm{R}x}\right].$ (A.1)
Hence, $h^{\prime}\geq 0$ and $h$ is increasing. The equality in (A.1) holds if and only if $\tilde{\mathbb{E}}\left[e^{-\frac{1}{2}\bm{R}x}\right]\bm{R}^{2}e^{-\frac{1}{2}\bm{R}x}=e^{-\frac{1}{2}\bm{R}x}\tilde{\mathbb{E}}\left[\bm{R}^{2}e^{-\frac{1}{2}\bm{R}x}\right]$, which is equivalent to $\bm{R}$ following a single-point distribution. Therefore, $h$ is strictly increasing unless $\bm{R}$ is a constant. ∎
@@
Let $\bm{R}$ be a non-negative random variable and let $r_{0}\triangleq\operatorname{essinf}\bm{R}$. Then $\lim\limits_{x\to\infty}h(x)=\frac{1}{r_{0}}$, where we adopt the convention $\frac{1}{0}=\infty$.
@@
where $Q(x)\triangleq\frac{\tilde{\mathbb{E}}[\bm{X}e^{-\frac{1}{2}\bm{X}x}]}{\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{X}x}]}$. By $\operatorname{essinf}\bm{X}=0$, we have $p_{\varepsilon}\triangleq\tilde{\mathbb{P}}(0\leq\bm{X}\leq\varepsilon)>0$ for any $\varepsilon>0$. Therefore,
$Q(x)\leq$ $\frac{\tilde{\mathbb{E}}[\bm{X}e^{-\frac{1}{2}\bm{X}x}\mathbb{I}_{\{\bm{X}\leq\varepsilon\}}+\tilde{\mathbb{E}}[\bm{X}e^{-\frac{1}{2}\bm{X}x}\mathbb{I}_{\{\bm{X}>\varepsilon\}}]}{\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{X}x}\mathbb{I}_{\{\bm{X}\leq\varepsilon\}}]}$
$\leq$ $\varepsilon+\frac{\tilde{\mathbb{E}}[\bm{X}e^{-\frac{1}{2}\bm{X}x}\mathbb{I}_{\{\bm{X}>\varepsilon\}}]}{p_{\varepsilon}e^{-\frac{1}{2}\varepsilon x}}=\varepsilon+\frac{\tilde{\mathbb{E}}[\bm{X}e^{-\frac{1}{2}(\bm{X}-\varepsilon)x}\mathbb{I}_{\{\bm{X}>\varepsilon\}}]}{p_{\varepsilon}}.$
Note that for any $x\geq 1$, $\bm{X}e^{-\frac{1}{2}(\bm{X}-\varepsilon)x}\mathbb{I}_{\{\bm{X}>\varepsilon\}}$ is dominated by $\bm{X}e^{-\frac{1}{2}(\bm{X}-\varepsilon)}\mathbb{I}_{\{\bm{X}>\varepsilon\}}\in L^{\infty}$. Hence, by the dominated convergence theorem, $\limsup\limits_{x\to\infty}Q(x)\leq\varepsilon$. Because $\varepsilon$ is arbitrary, we have $\lim\limits_{x\to\infty}Q(x)=0$. By (A.2), the convergence result for $h$ follows immediately. ∎
@@
The sequence $\{\tilde{\mathbb{E}}[\bm{R}_{n}]\}_{n\geq 1}$ is bounded because $\tilde{\mathbb{E}}[\bm{R}_{n}]\to\tilde{\mathbb{E}}[\bm{R}]$. Hence, the family of functions $\{\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{R}_{n}x}]\}_{n\geq 1}$ is equicontinuous on $\mathcal{K}$. Moreover, we have $\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{R}_{n}x}]\to\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{R}x}]$ for any $x\geq 0$, because $\bm{R}_{n}\xrightarrow{d}\bm{R}$ and the function $r\mapsto e^{-\frac{1}{2}rx}$ is bounded and continuous. It follows that $\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{R}_{n}x}]\to\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{R}x}]$ uniformly on $\mathcal{K}$; see, e.g., Rudin (1976, Exercise 7.16).
Now we show the uniform convergence of $\{\tilde{\mathbb{E}}[\bm{R}_{n}e^{-\frac{1}{2}\bm{R}_{n}x}]\}_{n\geq 1}$ on $\mathcal{K}$. By $\bm{R}_{n}\xrightarrow{d}\bm{R}$ and $\tilde{\mathbb{E}}[\bm{R}_{n}]\to\tilde{\mathbb{E}}[\bm{R}]<\infty$, we know that the family $\{\bm{R}_{n}\}_{n\geq 1}$ is uniformly integrable. Consequently, for any $\varepsilon>0$, there exists $M>0$ such that $F_{\bm{R}}$ is continuous at $M$, $\sup_{n}\tilde{\mathbb{E}}[\bm{R}_{n}\mathbb{I}_{\{\bm{R}_{n}>M\}}]<\frac{\varepsilon}{3}$, and $\tilde{\mathbb{E}}[\bm{R}\mathbb{I}_{\{\bm{R}>M\}}]<\frac{\varepsilon}{3}$. Let $f_{n}^{M}(x)\triangleq\tilde{\mathbb{E}}[\bm{R}_{n}e^{-\frac{1}{2}\bm{R}_{n}x}\mathbb{I}_{\{\bm{R}_{n}\leq M\}}]$. Then $f_{n}^{M}$ is continuously differentiable and its derivative is uniformly bounded by $\frac{1}{2}M^{2}$, which implies that the family $\{f_{n}^{M}\}_{n\geq 1}$ is equicontinuous on $\mathcal{K}$. Furthermore, because $\bm{R}_{n}\xrightarrow{d}\bm{R}$, we have the pointwise convergence $f_{n}^{M}(x)\to f^{M}(x)\triangleq\tilde{\mathbb{E}}[\bm{R}e^{-\frac{1}{2}\bm{R}x}\mathbb{I}_{\{\bm{R}\leq M\}}]$; see, e.g., Föllmer and Schied (2016, Theorem A.43). As a result, $f_{n}^{M}$ converges uniformly to $f^{M}$ on $\mathcal{K}$. Hence, for all sufficiently large $n$,
@@
$\left|\tilde{\mathbb{E}}[\bm{R}_{n}e^{-\frac{1}{2}\bm{R}_{n}x}]-\tilde{\mathbb{E}}[\bm{R}e^{-\frac{1}{2}\bm{R}x}]\right|\leq|f_{n}^{M}(x)-f^{M}(x)|+\tilde{\mathbb{E}}[\bm{R}_{n}\mathbb{I}_{\{\bm{R}_{n}>M\}}]+\tilde{\mathbb{E}}[\bm{R}\mathbb{I}_{\{\bm{R}>M\}}].$
@@
Finally, the function $\tilde{\mathbb{E}}[\bm{R}e^{-\frac{1}{2}\bm{R}x}]$ is bounded away from zero on $\mathcal{K}$, given that $\tilde{\mathbb{E}}[\bm{R}]>0$. Therefore, $\{h_{\bm{R}_{n}}\}_{n\geq 1}$ converges uniformly to $h_{\bm{R}}$ on $\mathcal{K}$. ∎
The reverse hazard rate order can be characterized via expectation ratios, as established by Capéraà (1988) (see also Shaked and Shanthikumar (2007, Theorem 1.B.50)).
@@
Let $\bm{R}_{1}$} and $\bm{R}_{2}$} be two random variables. Then ${\bm{R}_{1}}\succeq_{rh}{\bm{R}_{2}}$ if and only if
$\frac{\tilde{\mathbb{E}}\left[u({\bm{R}_{1}})w({\bm{R}_{1}})\right]}{\tilde{\mathbb{E}}[w({\bm{R}_{1}})]}\geq\frac{\tilde{\mathbb{E}}[u({\bm{R}_{2}})w({\bm{R}_{2}})]}{\tilde{\mathbb{E}}[w({\bm{R}_{2}})]}$
for all functions $u$ and $w$ such that the expectations exist, $w$ is positive and decreasing, and $u$ is increasing.
The next lemma shows that $h_{\bm{R}}$ is “decreasing” w.r.t. $\bm{R}$ (under the reverse hazard rate order of $\bm{R}$).
Lemma A.5.
Let $\bm{R}_{1}$ and $\bm{R}_{2}$ be two non-negative random variables with $F_{{\bm{R}_{1}}}\neq F_{{\bm{R}_{2}}}$. If $\bm{R}_{1}\succeq_{rh}\bm{R}_{2}$, then $h_{1}(x)<h_{2}(x)$ for all $x\geq 0$, where $h_{i}=h_{\bm{R}_{i}}$ ($i=1,2$).
@@
Fix $x\geq 0$. Let $w(\gamma)=\exp(-\frac{1}{2}\gamma x)$ and $u(\gamma)=\gamma$. Applying Lemma A.4 yields
$\frac{\tilde{\mathbb{E}}[{\bm{R}_{1}}w({\bm{R}_{1}})]}{\tilde{\mathbb{E}}[w({\bm{R}_{1}})]}\geq\frac{\tilde{\mathbb{E}}[{\bm{R}_{2}}w({\bm{R}_{2}})]}{\tilde{\mathbb{E}}[w({\bm{R}_{2}})]},$ (A.3)
which implies that $h_{1}(x)\leq h_{2}(x)$. To show the strict inequality, suppose on the contrary that the equality holds in (A.3). Let $\tilde{F}_{i}$ ($i=1,2$) be the probability distribution functions defined by $\mathrm{d}\tilde{F}_{i}(\gamma)=\frac{w(\gamma)\mathrm{d}F_{\bm{R}_{i}}(\gamma)}{\tilde{\mathbb{E}}[w(\bm{R}_{i})]}$. Then $\int_{0}^{\infty}\gamma\mathrm{d}\tilde{F}_{1}(\gamma)=\int_{0}^{\infty}\gamma\mathrm{d}\tilde{F}_{2}(\gamma)$. On the other hand, a combination of $\bm{R}_{1}\succeq_{rh}\bm{R}_{2}$, Lemma A.4, and the definition of $\tilde{F}_{i}$ yields $\tilde{F}_{1}\succeq_{1}\tilde{F}_{2}$. Therefore, $\tilde{F}_{1}=\tilde{F}_{2}$. By $w(\cdot)>0$, we have $F_{{\bm{R}_{1}}}=F_{{\bm{R}_{2}}}$, which contradicts the assumption that $F_{{\bm{R}_{1}}}\neq F_{{\bm{R}_{2}}}$. Thus, the inequality must be strict. ∎